### v1

# Iarrobino’s decomposition for self-dual modules

Abstract
We investigate self-dual modules and their Hilbert functions. We generalize Iarrobino’s decomposition to this setup and deduce consequences for their local Hilbert functions. We classify the local Hilbert functions for small degree modules. We prove a generalization of Kunte’s criterion for self-duality in terms of Macaulay’s inverse systems.
1 Introduction
1.1 Main results
In our work we introduce three main results: Iarrobino’s decomposition for self-dual modules, the generalization of Kunte’s criterion for self-duality in terms of Macaulay’s inverse systems and the classification of possible Hilbert functions of self-dual modules for degree less or equal to eight.
A. Iarrobino introduced Iarrobino’s decomposition for self-dual algebras [Iarrobino1994AssociatedGA, IARROBINO2021106496]. It is used in the deformation theory [Casnati2014IrreducibilityOT], the theory of symmetric matrices [Kosir], the classical algebraic geometry [Schreyer1999GeometryAA, ArzZai22], the theory of higher-dimensional Gorenstein algebras [ELIAS2017306, Jelisiejew2015ClassifyingLA]. The main claim in this result is that the local Hilbert function of a self-dual algebra can be decomposed as a sum of symmetric sequences. This result introduces criteria that the local Hilbert function of a self-dual algebra must satisfy.
F. Macaulay introduced the bound on the Hilbert functions of algebras [Macaulay]. This result was generalized to modules by H. Hulett [Hulett]. Iarrobino’s decomposition in combination with Macaulay Bound is a very useful tool in showing that a sequence is not a local Hilbert function of any self-dual algebra.
We generalize the result of A.Iarrobino to the case of self-dual modules. For our best knowledge this is the first time that Iarrobino’s decomposition is considered in terms of modules. Put informally, for a module $M$ of the socle degree $d$, the local Hilbert function $h_{M}$ is a sum
$h_{M}=\sum_{i=0}^{d}\Delta_{i},$
where $\Delta_{i}$ is symmetric around $\frac{d-i}{2}$. We now introduce the necessary definitions (see also Section 3).
Let $R$ be a local ring and let $\mathfrak{m}$ be its maximal ideal. Let $d$ be a maximal natural number $d$ such that $\mathfrak{m}^{d}M\neq 0$. By $\operatorname{gr}R$ we denote the associated graded ring of $R$, i.e.
$\operatorname{gr}R=\bigoplus_{k=0}^{d}\frac{\mathfrak{m}^{k}}{\mathfrak{m}^{k+1}}.$
By $\operatorname{gr}M$ we denote the associated graded module of the module $M$, i.e.
$\operatorname{gr}M=\bigoplus_{k=0}^{d}\frac{\mathfrak{m}^{k}M}{\mathfrak{m}^{k+1}M},$
which is a ($\operatorname{gr}R$)-module in a natural way. For an ideal $\mathfrak{a}$ in $R$ we denote by $(0:\mathfrak{a})_{M}$ the annihilator of $\mathfrak{a}$ in $M$, i.e. the set of all elements $x\in M$, such that $\mathfrak{a}x=0$. The annihilator $(0:\mathfrak{a})_{M}$ is a submodule of $M$.
Let us denote
$Q_{k,\>l}=\frac{\mathfrak{m}^{k}M\cap(0:\mathfrak{m}^{l})_{M}}{\mathfrak{m}^{k+1}M\cap(0:\mathfrak{m}^{l})_{M}+\mathfrak{m}^{k}M\cap(0:\mathfrak{m}^{l-1})_{M}}\text{ for all }0\leq k,\>l\leq d+1.$
Let us also define
[…]
We have $\operatorname{rank}_{\mathbb{K}}\>Q_{k,\>l}=\Delta_{d+1-(k+l)}(k)$.
We denote by $h_{M}$ the local Hilbert function of the module $M$, i.e.
$h_{M}(k)=\dim_{\mathbb{K}}\frac{\mathfrak{m}^{k}M}{\mathfrak{m}^{k+1}M}.$
The theorem is stated as follows:
Theorem 1.1 (Hilbert function decomposition).
If $R$ is a local ring and $M$ is a self-dual module, then the following statements hold true:
  1. 1.
[…]
  2. 2.
    For $k+l>d+1$ we have the identity $A_{k,\>l}=A_{k,\>l-1}$ and so it holds that $Q_{k,\>l}=0$.
  3. 3.
    It holds that $\Delta_{s}(t)=\Delta_{s}((d-s)-t)$. We say that $\Delta_{s}$ is symmetric with respect to $\frac{d-s}{2}$.
  4. 4.
    For $0\leq a\leq d$ let us denote $h_{a}(t)=\sum_{i=0}^{a}\Delta_{i}(t)$. Then $h_{a}$ is the Hilbert function of the quotient
    $\frac{\operatorname{gr}\>M}{\bigoplus_{i=0}^{a}C_{a}^{i}},$
    where $C_{a}^{i}\subseteq\mathfrak{m}^{i}M/\mathfrak{m}^{i+1}M$ is the image of $A_{i,d-(i+a)}$.
  5. 5.
    The local Hilbert function $h_{M}$ satisfies $h_{M}(t)=0$ for $t>d$ and
    $h_{M}(t)=\sum_{i=0}^{d-t}\Delta_{i}(t)\text{ for }0\leq t\leq d.$
As examples of applications of this theorem we show two claims that we also note in Section 5.
Example 1.2.
  • If $h_{M}$ is a Hilbert function of a self-dual module $M$ and $k$ is the largest natural number such that $h_{M}(k)>0$, then $h_{M}(0)\geq h_{M}(k)$.
    Proof.
    Let us consider Iarrobino’s decomposition of $h_{M}$ – first row must be symmetric (from Theorem 1.1 Point 4) and must have $h_{M}(k)$ on the last position (since next rows are strictly shorter), so it also must have $h_{M}(k)$ on the first position. Since $h_{M}(0)$ is a sum of values on the first positions in all rows, we have $h_{M}(0)\geq h_{M}(k)$. ∎
  • Let $h_{M}$ be a Hilbert function of a self-dual module $M$ and let $k$ be the largest natural number such that $h_{M}(k)>0$. Then $\sum_{i=0}^{\lfloor{\frac{k-1}{2}}\rfloor}h_{M}(i)\geq\sum_{i=\lceil{\frac{k+1}{2}}\rceil}^{k}h_{M}(i)$.
    Proof.
    The claim states that the sum of the first half of the sequence is greater or equal to the sum of the second half. This follows from Iarrobino’s decomposition (Theorem 1.1) – the sequence is a sum of symmetric parts and each of these parts contributes to the first half at least as much as it does to the second half. ∎
J. Jelisiejew and K. Šivic introduced the concept of apolarity for modules, which is a very useful tool in generating examples of modules with given Hilbert function [Jelisiejew2022ComponentsAS]. M. Kunte introduced a sufficient and necessary condition for a module to be a graded self-dual module [Kunte2008GorensteinMO]. We use the notion of apolarity to prove a generalization of Kunte’s criterion to the non-graded case. However, we show the criterion with only a sufficient part, while the criterion by Kunte for the graded case was also necessary. The theorem is stated as follows: Let $n$ be a positive natural number and let us denote $S^{*}=\mathbb{K}[x_{1},\>x_{2},\>\ldots,\>x_{n}]$. Let us also fix a positive natural number $r$. We denote by $F^{*}$ the free $S^{*}$-module $\bigoplus_{k=0}^{r}S^{*}e_{k}^{*}$. Note that each $f_{k}\in F^{*}$ is encoded by the direct sum decomposition as $\sum_{i=0}^{r}f_{ki}e_{i}^{*}$, where $f_{ki}\in S^{*}$. Thus, the tuple $f_{1},\>f_{2},\ldots,f_{l}$ can be viewed as a matrix $N=[f_{ki}]\in\mathbb{M}_{l\times r}(S^{*})$. We call this matrix the encoding matrix of the module apolar to $(f_{1},\>f_{2},\ldots,f_{l})$.
Theorem 1.3 (Kunte’s criterion).
If the square matrix $N=[f_{ki}]\in\mathbb{M}_{r\times r}(S^{*})$ is symmetric, then the module apolar to $(f_{k})_{k=1}^{r}$ is self-dual, where $f_{k}\in F^{*}$.
A definition of the apolarity notion can be found in Section 4. The Kunte’s criterion is a quite simple condition, as we only need to verify if the proper matrix is symmetric. This criterion, combined with apolarity for modules, allows us to easily generate examples of self-dual modules for a given Hilbert function. The special case of this theorem is $r=1$. Then $M$ becomes an algebra. Since the $1\times 1$ matrix is trivially symmetric, it implies that every apolar algebra is self-dual.
Example 1.4.
Let us take as the encoding matrix the diagonal matrix
$\begin{bmatrix}x^{2}&0\\ 0&x\\ \end{bmatrix}.$
The obtained self-dual module has the local Hilbert function $(2,2,1)$. The method for computing the local Hilbert function from the encoding matrix is described in Section 4 (see Example 4.10). The reader might observe that it corresponds to the direct sum decomposition $(1,1,1)+(1,1)$ (see Section 5).
We show other applications of this theorem in Example 4.10 and in the classification of the local Hilbert function $(2,2,3,1)$ (see Section 5).
B. Poonen classified the local Hilbert functions and the types of isomorphisms for self-dual algebras of degree less or equal to six [Poonen]. The corresponding problem for modules is mostly open. In our work we classify possible local Hilbert functions of self-dual modules with a small degree – we classify all possible Hilbert functions for a degree less or equal to $8$.
Proposition 1.5.
The number of possible local Hilbert functions of self-dual modules for degree $m=1,2\ldots 8$ is exactly the following:
Table 1: Number of possible local Hilbert functions of self-dual modules
Degree Number of possible local Hilbert functions
$m=1$ $1$
$m=2$ $2$
$m=3$ $3$
$m=4$ $6$
$m=5$ $9$
$m=6$ $16$
$m=7$ $24$
$m=8$ $38$
The whole classification of possible local Hilbert functions of self-dual modules can be found in Section 5. This classification is interesting, as we show that some unexpected local Hilbert functions are possible to obtain from self-dual modules, for example the function $(2,2,3,1)$.
1.2 Self-dual algebras and modules
Self-dual algebras (also known as Gorenstein algebras or commutative Frobenius algebras) are useful tools in algebraic complexity theory. Let us consider a tensor $t\in A\otimes B\otimes C$, where $A,\>B,\>C\cong\mathbb{C}^{m}$. The tensor $t$ may be considered as a linear map $A^{*}\rightarrow B\otimes C$. We say that $t$ is $1_{A}$-generic if there exists an $\alpha$ such that $t(\alpha):B^{*}\to C$ has full rank. We define analogically $1_{B}$- and $1_{C}$-genericity. We say that a tensor is $1$-generic, if it is $1_{A}$-, $1_{B}$- and $1_{C}$-generic. Let us assume that $t$ is $1_{A}$-generic. Let $\alpha$ be such an element that $t(\alpha)$ is invertible. Let us consider the space $t(A^{*})\cdot t(\alpha)^{-1}$. We say that $t$ satisfies the Strassen’s equations, if the space $t(A^{*})\cdot t(\alpha)^{-1}$ is abelian, i.e. if it consists of commuting matrices. More about Strassen’s equations one can read in [LaMM]. A tensor is a structure tensor of an algebra $A$ if it corresponds to the multiplication map $A\times A\rightarrow A$. The following result connects $1$-generic tensors satisfying Strassen’s equations and self-dual algebras:
Proposition 1.6.
([Jelisiejew2022ConciseTO, Summary 2.5]) Let us assume that we have $A,\>B,\>C\cong\mathbb{C}^{m}$. A tensor $t\in A\otimes B\otimes C$ is $1$-generic and satisfies Strassen’s equations if and only if it is a structure tensor of a self-dual algebra.
The self-dual algebras are also a useful tool for investigating secant varieties, see [MR3121848]. They are subjects of intensive research, see [MR2533305, MR3404648].
In this work we investigate self-dual modules (see Definition 2.1) as they are also important in algebraic complexity theory. Deriving algebraic structure on tensors allows to analyze tensors of matrix multiplication and obtain bounds on $\omega$ [strassen1969gaussian, Alman2021ARL, blser_et_al:LIPIcs:2020:12686, Wojtala2022], which is one of the most important topics in theoretical computer science. M. Wojtala introduced the notion of structure tensors of modules [Wojtala2022]. The following result connects $1_{A}$-generic tensors satisfying Strassen’s equations and modules:
Proposition 1.7.
([Jelisiejew2022ConciseTO, Summary 2.5], [LaMM, Lemma 2.6], [Wojtala2022, Proof of Corollary 2]) Let us assume that we have $A,\>B,\>C\cong\mathbb{C}^{m}$. A tensor $t\in A\otimes B\otimes C$ is $1_{A}$-generic and satisfies Strassen’s equations if and only if it is a structure tensor of a module.
For researchers working with tensors, it is usual to impose symmetry conditions. A tensor $t\in A\otimes A\otimes A$ is symmetric if it is invariant under the permutation of coordinates. By Proposition 1.6, one cannot hope for symmetric $1_{A}$-generic tensors, which do not come from algebras. However, imposing the partial symmetry is actually connected to coming from self-dual modules. This connection is shown in the following proposition and can be viewed as an analogical result to Proposition 1.6 and Proposition 1.7.
Proposition 1.8.
Let us assume that we have $A,\>B\cong\mathbb{C}^{m}$. If a tensor $t\in A\otimes S^{2}(B)$ is $1_{A}$-generic and satisfies Strassen’s equations, then it is a structure tensor of a self-dual module.
Proof.
Proposition 1.7 implies that $t$ is a structure tensor of some module $M$. Now let us observe that by symmetry of matrices from $t(A^{*})$ and a reasoning from [Jelisiejew2022ComponentsAS, Subsection 3.5] we have that $M$ is isomorphic to $M^{*}$ as modules. ∎
Acknowledgements
During the preparation of this publication, the author was part of the Szkoła Orłów program. The publication was created under the supervision of Joachim Jelisiejew, whose support and help were priceless.
2 Preliminaries
In this section we recall facts that will be useful for our purposes.
Let $\mathbb{K}$ denote an algebraically closed field. Let also $R$ denote a ring that is a finite-dimensional $\mathbb{K}$-vector space and let $M$ denote an $R$-module that is a finite-dimensional $\mathbb{K}$-vector space. The degree of such a module $M$ is defined to be the dimension of $M$ as a $\mathbb{K}$-vector space.
On a dual vector space $M^{*}$ we introduce an $R$-module structure by defining
$r\cdot\phi(m)=\phi(rm).$
Definition 2.1.
A module $M$ is self-dual if it is isomorphic with $M^{*}$ as an $R$-module.
A case of interest is when $R$ is a local ring and $M$ is a self-dual module, although some results are shown without those assumptions.
Definition 2.2.
For an ideal $\mathfrak{a}$ in $R$ we denote $\mathfrak{a}^{z}=(1)$ for any negative integer $z$.
2.1 Modules and submodules
Definition 2.3.
For a submodule $K$ of $M$ we define $K^{\perp}$ by the formula $K^{\perp}=\left\{\phi\in M^{*}:\phi(K)=0\right\}$.
Lemma 2.4.
For a module $K$ we have that $K^{\perp}$ is a submodule of $M^{*}$.
Proof.
If $\phi_{1},\phi_{2}\in K^{\perp}$, then $\phi_{1}(K)=0,\phi_{2}(K)=0$, so $(\phi_{1}+\phi_{2})(K)=\phi_{1}(K)+\phi_{2}(K)=0$, so $\phi_{1}+\phi_{2}\in K^{\perp}$. Also if $\phi\in K^{\perp}$, then for every $r\in R$ we have that $rK\subseteq K$ (since $K$ is a submodule of $M$) and thus $(r\cdot\phi)(K)=\phi(rK)=0$, so $r\cdot\phi\in K$. ∎
Lemma 2.5.
  1. 1.
    Let $K$ be a submodule of $M$. Then $\dim_{\mathbb{K}}K+\dim_{\mathbb{K}}K^{\perp}=\dim_{\mathbb{K}}M$.
  2. 2.
    Let $K$ be a submodule of $M$. Then $(K^{\perp})^{\perp}\cong K$.
  3. 3.
    Let $K,\>L$ be submodules of $M$. Then we have $(K+L)^{\perp}=K^{\perp}\cap L^{\perp}$.
  4. 4.
    Let $K,\>L$ be submodules of $M$. Then $(K\cap L)^{\perp}=K^{\perp}+L^{\perp}$.
Proof.
Left to the reader. ∎
Lemma 2.6.
We have a canonical isomorphism $M\cong M^{**}$ given by the double dual map.
Proof.
Let $\psi:M\rightarrow M^{**}$ be the double dual map, i.e. $\psi(m)(\phi)=\phi(m)$. Then $\psi$ is an isomorphism of $\mathbb{K}$-linear spaces, we need to show that it is also an isomorphism of $R$-modules. Indeed, we have $s\cdot\psi(m)(\phi)=\psi(m)(s\phi)=(s\phi)(m)=\phi(sm)=\psi(s\cdot m)(\phi)$. ∎
Lemma 2.7.
Let $I$ be a submodule of $M$. Then we have a canonical isomorphism $I^{*}\cong\nicefrac{{M^{*}}}{{I^{\perp}}}$ induced by the canonical surjection.
Proof.
Let $\pi:M^{*}\rightarrow I^{*}$ be the canonical surjection. Then $\ker\pi=\left\{\phi\in M^{*}:\phi(I)=0\right\}=I^{\perp}$. So by the isomorphism theorem we have a canonical isomorphism $I^{*}\cong\nicefrac{{M^{*}}}{{\ker\pi}}=\nicefrac{{M^{*}}}{{I^{\perp}}}$. ∎
Lemma 2.8.
Let $J$ be a submodule of $I$ and $I$ be a submodule of $M$. Then we have a canonical exact sequence $0\rightarrow(\nicefrac{{I}}{{J}})^{*}\rightarrow I^{*}\rightarrow J^{*}\rightarrow 0$.
Proof.
We obtain the canonical injection $(\nicefrac{{I}}{{J}})^{*}\rightarrow I^{*}$ by lifting functionals from $(\nicefrac{{I}}{{J}})^{*}$ to functionals from $I^{*}$, i.e. functional $\phi$ goes to functional $\overline{\phi}$, such that for $i\in I$, $i\equiv\overline{i}\mod J$ we have $\overline{\phi}(i)=\phi(\overline{i})$. The image of this injection consists of all functionals that are zero on $J$. The canonical surjection $I^{*}\rightarrow J^{*}$ is obtained by restricting functionals on $I$ to $J$, its kernel also consists of all functionals that are zero on $J$, so we constructed the desired canonical exact sequence. ∎
Lemma 2.9.
Let $J$ be a submodule of $I$ and $I$ be a submodule of $M$. Then $(\nicefrac{{I}}{{J}})^{*}\cong\nicefrac{{J^{\perp}}}{{I^{\perp}}}$.
Proof.
Left to the reader. ∎
Lemma 2.10.
Let $M_{1}\>M_{2}$ be $R$-modules that are finite-dimensional $\mathbb{K}$-vector spaces. Let us assume that there exists a non-degenerate pairing $\langle\>,\>\rangle:M_{1}\times M_{2}\rightarrow\mathbb{K}$ such that for every $r\in R,\>a_{1}\in M_{1},\>a_{2}\in M_{2}$, we have $\langle ra_{1},\>a_{2}\rangle=\langle a_{1},\>ra_{2}\rangle$. Then $M_{1}$ and $M_{2}$ are dual as $R$-modules.
Proof.
Left to the reader. ∎
Definition 2.11.
For an ideal $\mathfrak{a}$ in $R$ we denote by $(0:\mathfrak{a})_{M}$ the annihilator of $\mathfrak{a}$ in $M$, i.e. the set of all elements $x\in M$, such that $\mathfrak{a}x=0$. It is easy to check that $(0:\mathfrak{a})_{M}$ is a submodule of $M$.
Lemma 2.12.
Let $\mathfrak{a}$ be an ideal in $R$. If $M$ is self-dual, then we have the following isomorphism of $R$-modules: $(\mathfrak{a}M)^{\perp}\cong(0:\mathfrak{a})_{M}$.
Proof.
We have $(\mathfrak{a}M)^{\perp}=\left\{\phi\in M^{*}:\phi(\mathfrak{a}M)=0\right\}=\left\{\phi\in M^{*}:\forall_{s\in\mathfrak{a}}{s\phi=0}\right\}\cong\left\{l\in M:\forall_{s\in\mathfrak{a}}sl=0\right\}=(0:\mathfrak{a})_{M}$, where the isomorphism follows from $M\cong M^{*}$. ∎
2.2 Local rings
In this subsection we assume that $R$ is a local ring and $\mathfrak{m}$ is its only maximal ideal.
Corollary 2.13.
If $M$ is self-dual, then we have the following isomorphisms of $R$-modules: $\left(\mathfrak{m}^{k}M\right)^{\perp}\cong(0:\mathfrak{m}^{k}M)_{M}$, $\left((0:\mathfrak{m}^{k}M)_{M}\right)^{\perp}\cong\mathfrak{m}^{k}M$.
Proof.
The first statement is just Lemma 2.12 for the ideal $\mathfrak{m}^{k}$. The second part is obtained from the first one by taking the perpendicular submodule on both sides and using Lemma 2.5 Point 2. ∎
Remark 2.14.
There exists the largest natural number $d$ such that $\mathfrak{m}^{d}M\neq 0$.
2.3 Modularity and filtrations
Remark 2.15.
Let $A,\>B,\>C$ be submodules of $M$ and let us assume that $A$ is a submodule of $C$. Then $(A+B)\cap C=B\cap C+A$.
Corollary 2.16.
Let $A_{1},\>A_{2},\>B_{1},\>B_{2}$ be submodules of $M$ and let us also assume that $A_{1}$ is a submodule of $B_{1}$ and $B_{2}$ is a submodule of $A_{2}$. Then $(A_{1}+A_{2})\cap(B_{1}+B_{2})=A_{1}+B_{2}+A_{2}\cap B_{1}$.
Proof.
First we use Remark 2.15 for $A=A_{1},\>B=A_{2},\>C=(B_{1}+B_{2})$ and obtain $(A_{1}+A_{2})\cap(B_{1}+B_{2})=A_{1}+A_{2}\cap(B_{1}+B_{2})$. Then we use Remark 2.15 for $A=B_{2},\>B=B_{1},\>C=A_{2}$ and obtain $A_{2}\cap(B_{1}+B_{2})=B_{2}+B_{1}\cap A_{2}$. ∎
Remark 2.17.
[…]
2.4 Associated graded ring and module
In this subsection we assume that $R$ is a local ring and $\mathfrak{m}$ is its only maximal ideal. We denote by $d$ the largest natural number such that $\mathfrak{m}^{d}\neq 0$, as in Remark 2.14.
Definition 2.18.
By $\operatorname{gr}R$ we denote the associated graded ring of $R$, i.e.
$\operatorname{gr}R=\bigoplus_{k=0}^{d}\frac{\mathfrak{m}^{k}}{\mathfrak{m}^{k+1}}.$
The multiplication in this ring is defined as follows: for $a_{k}\in\mathfrak{m}^{k}$ let $\overline{a_{k}}$ denote the image of $a_{k}$ in $\frac{\mathfrak{m}^{k}}{\mathfrak{m}^{k+1}}$; then $\overline{a_{i}}\cdot\overline{a_{j}}$ is defined as $\overline{a_{i}\cdot a_{j}}$, i.e. it is the image of $a_{i}\cdot a_{j}$ in $\frac{\mathfrak{m}^{i+j}}{\mathfrak{m}^{i+j+1}}$.
By $\operatorname{gr}M$ we denote the associated graded module of the module $M$, i.e.
$\operatorname{gr}M=\bigoplus_{k=0}^{d}\frac{\mathfrak{m}^{k}M}{\mathfrak{m}^{k+1}M},$
which is a ($\operatorname{gr}R$)-module in a natural way.
Definition 2.19.
We denote by $h_{M}$ the local Hilbert function of the module $M$, i.e.
$h_{M}(k)=\dim_{\mathbb{K}}\frac{\mathfrak{m}^{k}M}{\mathfrak{m}^{k+1}M}.$
Let us denote $\mathbb{n}=\bigoplus_{k>0}\frac{\mathbb{m}^{k}}{\mathbb{m}^{k+1}}\subseteq\operatorname{gr}R$. Clearly $\mathbb{n}$ is an ideal in $\operatorname{gr}R$. The following lemma shows that is the unique maximal ideal in $\operatorname{gr}R$.
Lemma 2.20.
The associated graded ring $\operatorname{gr}R$ is a local ring and $\mathbb{n}$ is its unique maximal ideal.
Proof.
We have
$\frac{\operatorname{gr}R}{\mathbb{n}}=\frac{\operatorname{gr}R}{\bigoplus_{k>0}\frac{\mathbb{m}^{k}}{\mathbb{m}^{k+1}}}\cong\frac{R}{\mathbb{m}}.$
Since $\frac{R}{\mathbb{m}}$ is a field, we obtain that $\mathbb{n}$ is a maximal ideal. Moreover, $\mathbb{n}^{d+1}=(\bigoplus_{k>0}\frac{\mathbb{m}^{k}}{\mathbb{m}^{k+1}})^{d+1}=0$, so since maximal ideals are prime, every maximal ideal must contain $\mathbb{n}$. It follows that $\mathbb{n}$ is the only maximal ideal in $\operatorname{gr}R$. ∎
Lemma 2.21.
It holds that $\mathbb{n}^{l}=\bigoplus_{k\geq l}\frac{\mathbb{m}^{k}}{\mathbb{m}^{k+1}}$.
Proof.
Clearly $\mathbb{n}^{l}=(\bigoplus_{k\geq 1}\frac{\mathbb{m}^{k}}{\mathbb{m}^{k+1}})^{l}\subseteq\bigoplus_{k\geq l}\frac{\mathbb{m}^{k}}{\mathbb{m}^{k+1}}$. To show the opposite inclusion, let us fix an element $a\in\frac{\mathbb{m}^{l+c}}{\mathbb{m}^{l+c+1}}$, where $c\geq 0$. We can write $a$ as a finite sum $\sum_{j}a_{1,\>j}\cdot a_{2,\>j}\cdot\ldots\cdot a_{l+c,\>j}+\mathbb{m}^{l+c+1}$, where $a_{i,\>j}\in\mathbb{m}$. Then we have $a=\sum_{j}a_{1,\>j}\cdot a_{2,\>j}\cdot\ldots\cdot a_{l+c,\>j}+\mathbb{m}^{l+c+1}=\sum_{j}(a_{1,\>j}+\mathbb{m}^{2})\cdot(a_{2,\>j}+\mathbb{m}^{2})\cdot\ldots\cdot(a_{l+c,\>j}+\mathbb{m}^{2})\in(\frac{\mathbb{m}}{\mathbb{m}^{2}})^{l+c}\subseteq(\bigoplus_{k\geq 1}\frac{\mathbb{m}^{k}}{\mathbb{m}^{k+1}})^{l+c}\subseteq(\bigoplus_{k\geq 1}\frac{\mathbb{m}^{k}}{\mathbb{m}^{k+1}})^{l}$. ∎
Corollary 2.22.
Let $C=\bigoplus_{k=0}^{d}\frac{C_{k}+\mathbb{m}^{k+1}M}{\mathbb{m}^{k+1}M}$ be a graded submodule of $\operatorname{gr}M$, where $C_{k}\subseteq\mathbb{m}^{k}M$ is an $R$-submodule. Then the local Hilbert function $h_{q}$ of the quotient $\frac{\operatorname{gr}M}{C}$ satisfies $h_{q}(k)=\dim_{\mathbb{K}}\frac{\mathbb{m}^{k}M}{\mathbb{m}^{k+1}M+C_{k}}$ and the local Hilbert function $h_{C}$ of the module $C$ satisfies $h_{C}(k)=\dim_{\mathbb{K}}\frac{C_{k}+\mathbb{m}^{k+1}M}{\mathbb{m}^{k+1}M}$.
Proof.
by Lemma 2.21 it follows that
$\frac{\mathbb{n}^{l}}{\mathbb{n}^{l+1}}=\frac{\bigoplus_{k\geq l}\frac{\mathbb{m}^{k}}{\mathbb{m}^{k+1}}}{\bigoplus_{k\geq l+1}\frac{\mathbb{m}^{k}}{\mathbb{m}^{k+1}}}=\frac{\mathbb{m}^{l}}{\mathbb{m}^{l+1}}.$
So the direct sum decompositions of $\operatorname{gr}M$ coming from $(\frac{\mathbb{n}^{l}}{\mathbb{n}^{l+1}})_{l=0}^{d}$ and $(\frac{\mathbb{m}^{l}}{\mathbb{m}^{l+1}})_{l=0}^{d}$ are identical, so values of the Hilbert function can be obtained from the latter one, which we wanted to show. ∎
The following remark states that the associated graded ring $\operatorname{gr}R$ satisfies the same assumptions as $R$ and $\operatorname{gr}M$ satisfies the same assumptions as $M$ (we do not assume that $M$ is self-dual here). It implies that we can apply Macaulay’s Bound to $\operatorname{gr}M$. Note that $\operatorname{gr}M$ in general might not be self-dual, so we cannot apply Iarrobino’s decomposition to $\operatorname{gr}M$.
Remark 2.23.
If $R$ is a local ring, then the associated graded ring $\operatorname{gr}R$ is a local $\mathbb{K}$-algebra and is a finite-dimensional $\mathbb{K}$-vector space. The associated graded module $\operatorname{gr}M$ is a finite-dimensional $\mathbb{K}$-vector space.
2.5 Macaulay’s Bound
In this subsection we assume that $R$ is a local ring. We recall now a criterion (thanks to Macaulay) that the local Hilbert functions of modules have to satisfy. Unlike Iarrobino’s decomposition, this criterion is also valid for modules that are not self-dual. Before we recall the aforesaid Macaulay’s Bound, we need to introduce one representation system.
Definition 2.24.
Let $a,\>s$ be positive integers. Then there exist uniquely determined positive integers $a_{s}>a_{s-1}>a_{s-2}>\ldots>a_{k}\geq k>0$, such that we have the equality
$a=\binom{a_{s}}{s}+\binom{a_{s-1}}{s-1}+\binom{a_{s-2}}{s-2}+\ldots+\binom{a_{k}}{k}.$
We denote
$a^{\langle s\rangle}=\binom{a_{s}+1}{s+1}+\binom{a_{s-1}+1}{s}+\binom{a_{s-2}+1}{s-1}+\ldots+\binom{a_{k}+1}{k+1}.$
Theorem 2.25 (Macaulay’s Bound, [Hulett]).
We have the following bound on the local Hilbert function:
$h_{M}(r+1)\leq h_{M}(r)^{\langle r\rangle}$
for $r\geq 1$. The bound holds true for $M$ that are not self-dual modules.
Corollary 2.26.
Let us assume that it holds $h_{M}(r)\leq r$. Then we have $h_{M}(r+1)\leq h_{M}(r)$.
Proof.
If $r=0$, then the assumption states that $h_{M}(0)=0$, so $M=0$ and the claim holds true. Now let us assume that $r\geq 1$. Since we have $h_{M}(r)\leq r$, the representation of $h_{M}(r)$ is $\binom{r}{r}+\binom{r-1}{r-1}+\binom{r-2}{r-2}+\ldots+\binom{r-h_{M}(r)+1}{r-h_{M}(r)+1}$. Thus $h_{M}(r)^{\langle r\rangle}=h_{M}(r)$ and by Macaulay’s Bound 2.25 we obtain $h_{M}(r+1)\leq h_{M}(r)^{\langle r\rangle}=h_{M}(r)$. ∎
Corollary 2.27.
Let us assume that it holds $h_{M}(r)\leq r+1$ and $r\geq 1$. Then we have $h_{M}(r+1)\leq r+2$.
Proof.
If $h_{M}(r)\leq r$, then from Corollary 2.26 we obtain the claim. If $h_{M}(r)=r+1$, then the representation of $h_{M}(r)$ is $\binom{r+1}{r}$ and thus $h_{M}(r)^{\langle r\rangle}=r+2$. So by Macaulay’s Bound 2.25 we obtain $h_{M}(r+1)\leq h_{M}(r)^{\langle r\rangle}=r+2$. ∎
3 Iarrobino’s decomposition
In this section we introduce a new result – Iarrobino’s decomposition for modules. Iarrobino’s decomposition was known for algebras, but as far as we know, this is the first time when Iarrobino’s decomposition has been considered for modules. With properly defined objects, the generalization from algebras to modules is direct, but there are some subtle differences, see Example 3.3 We assume in this section that $R$ is a local ring and $\mathfrak{m}$ is its only maximal ideal. Additionally, we assume that $M$ is self-dual, i.e. $M$ and $M^{*}$ are isomorphic as $R$-modules.
Proposition 3.1.
If $R$ is a local ring and $M$ is a self-dual module, then we have the following isomorphism of $\mathbb{K}$-vector spaces
$\left(\frac{\mathfrak{m}^{i}M\cap(0:\mathfrak{m}^{j})_{M}}{\mathfrak{m}^{i+1}M\cap(0:\mathfrak{m}^{j})_{M}+\mathfrak{m}^{i}M\cap(0:\mathfrak{m}^{j-1})_{M}}\right)^{*}\cong\frac{(0:\mathfrak{m}^{i+1})_{M}\cap\mathfrak{m}^{j-1}M}{(0:\mathfrak{m}^{i})_{M}\cap\mathfrak{m}^{j-1}M+(0:\mathfrak{m}^{i+1})_{M}\cap\mathfrak{m}^{j}M}.$
Proof.
Using Lemma 2.9, Lemma 2.5 Point 3 and Lemma 2.5 Point 4 respectively, we obtain
$\left(\frac{\mathfrak{m}^{i}M\cap(0:\mathfrak{m}^{j})_{M}}{\mathfrak{m}^{i+1}M\cap(0:\mathfrak{m}^{j})_{M}+\mathfrak{m}^{i}M\cap(0:\mathfrak{m}^{j-1})_{M}}\right)^{*}\cong\frac{\left(\mathfrak{m}^{i+1}M\cap(0:\mathfrak{m}^{j})_{M}+\mathfrak{m}^{i}M\cap(0:\mathfrak{m}^{j-1})_{M}\right)^{\perp}}{\left(\mathfrak{m}^{i}M\cap(0:\mathfrak{m}^{j})_{M}\right)^{\perp}}\cong$ (1)
$\cong\frac{\left(\mathfrak{m}^{i+1}M\cap(0:\mathfrak{m}^{j})_{M}\right)^{\perp}\cap\left(\mathfrak{m}^{i}M\cap(0:\mathfrak{m}^{j-1})_{M}\right)^{\perp}}{\left(\mathfrak{m}^{i}M\cap(0:\mathfrak{m}^{j})_{M}\right)^{\perp}}\cong$
$\cong\frac{\left(\left(\mathfrak{m}^{i+1}M\right)^{\perp}+\left((0:\mathfrak{m}^{j})_{M}\right)^{\perp}\right)\cap\left(\left(\mathfrak{m}^{i}M\right)^{\perp}+\left((0:\mathfrak{m}^{j-1})_{M}\right)^{\perp}\right)}{\left(\mathfrak{m}^{i}M\right)^{\perp}+\left((0:\mathfrak{m}^{j})_{M}\right)^{\perp}}.$
Now from Corollary 2.13 we have
$\frac{\left(\left(\mathfrak{m}^{i+1}M\right)^{\perp}+\left((0:\mathfrak{m}^{j})_{M}\right)^{\perp}\right)\cap\left(\left(\mathfrak{m}^{i}M\right)^{\perp}+\left((0:\mathfrak{m}^{j-1})_{M}\right)^{\perp}\right)}{\left(\mathfrak{m}^{i}M\right)^{\perp}+\left((0:\mathfrak{m}^{j})_{M}\right)^{\perp}}\cong$ (2)
$\cong\frac{\left((0:\mathfrak{m}^{i+1})_{M}+\mathfrak{m}^{j}M\right)\cap\left((0:\mathfrak{m}^{i})_{M}+\mathfrak{m}^{j-1}M\right)}{(0:\mathfrak{m}^{i})_{M}+\mathfrak{m}^{j}M}.$
From Corollary 2.16 we obtain
$\frac{\left((0:\mathfrak{m}^{i+1})_{M}+\mathfrak{m}^{j}M\right)\cap\left((0:\mathfrak{m}^{i})_{M}+\mathfrak{m}^{j-1}M\right)}{(0:\mathfrak{m}^{i})_{M}+\mathfrak{m}^{j}M}\cong$ (3)
$\cong\frac{(0:\mathfrak{m}^{i})_{M}+\mathfrak{m}^{j}M+(0:\mathfrak{m}^{i+1})_{M}\cap\mathfrak{m}^{j-1}M}{(0:\mathfrak{m}^{i})_{M}+\mathfrak{m}^{j}M}.$
Now we can rewrite
$\frac{(0:\mathfrak{m}^{i})_{M}+\mathfrak{m}^{j}M+(0:\mathfrak{m}^{i+1})_{M}\cap\mathfrak{m}^{j-1}M}{(0:\mathfrak{m}^{i})_{M}+\mathfrak{m}^{j}M}\cong$ (4)
$\cong\frac{(0:\mathfrak{m}^{i+1})_{M}\cap\mathfrak{m}^{j-1}M}{\left((0:\mathfrak{m}^{i})_{M}+\mathfrak{m}^{j}M\right)\cap\left((0:\mathfrak{m}^{i+1})_{M}\cap\mathfrak{m}^{j-1}M\right)}\cong$
$\cong\frac{(0:\mathfrak{m}^{i+1})_{M}\cap\mathfrak{m}^{j-1}M}{(0:\mathfrak{m}^{i})_{M}\cap\mathfrak{m}^{j-1}M+(0:\mathfrak{m}^{i+1})_{M}\cap\mathfrak{m}^{j}M}.$
Let $d$ be a maximal natural number $d$ such that $\mathfrak{m}^{d}M\neq 0$ ($d$ exists as we observe in Remark 2.14). Let us denote
[…]